OpenStudy (zmtgaminghd):

What value(s) of x will make each equation below true? 2x-6=3x+1-x-7

Vocaloid (vocaloid):

well, let's start by combining like terms just focus on the right side of the equation for now: 3x + 1 - x -7 let's combine the x's together 3x - x = ?

OpenStudy (zmtgaminghd):

2?

OpenStudy (anonymous):

2x-6 = 3x+1-x-7 2x-6 = 3x-x+1-7 2x-6 = 2x -6 2x -2x -6+6 = 2x - 2x +6 -6 0 = 0 there you go, hope this help

OpenStudy (anonymous):

just re arrange the number like Volcaloid do up their and start doing subtrach/add if you see (like above), I normally switch numbers but I think just write this out like this should be easier to understand to you :)